# On the dynamics of leapfrogging vortex rings

### v1

@@ Appendix C No collision in finite time
Although this fact is probably well-known, we provide the proof here since we did not find it in the literature.
As observed in [18], the system (C.1) is a Hamiltonian system with Hamiltonian given by
$H_{0}=\frac{1}{2}\sum_{i\neq j}\ln|q_{i}(t)-q_{j}(t)|+\frac{1}{4(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)^{2}.$
@@
$I(t)=\frac{1}{2}\sum_{i=1}^{N}|q_{i}(t)|^{2}$
satisfies $I(t)\leqslant I(0)e^{\gamma t/4\pi(r^{\ast})^{2}}$.
Indeed, we compute
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}I(t)&=\frac{\gamma}{2\pi}\sum_{i\neq j}q_{i}(t)\cdot\frac{(q_{i}(t)-q_{j}(t))^{\perp}}{|q_{i}(t)-q_{j}(t)|^{2}}-\frac{\gamma}{4\pi(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)q_{i,z}(t)\\ &=-\frac{\gamma}{4\pi(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)q_{i,z}(t)\leqslant\frac{\gamma}{4\pi(r^{\ast})^{2}}I(t),\end{split}$
where we have exchanged $i$ and $j$ in the first sum.
As a consequence, we get by standard arguments
$\begin{split}\sum_{i\neq j||q_{i}-q_{j}|<1}|\ln|q_{i}(t)-q_{j}(t)||&=\sum_{i\neq j||q_{i}-q_{j}|>1}\ln|q_{i}(t)-q_{j}(t)|-2H_{0}+\frac{1}{2(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)^{2}\\ &\leqslant 2\sqrt{2N}\sqrt{I(t)}-2H_{0}+\frac{1}{(r^{\ast})^{2}}I(t).\end{split}$
for all $t\in[0,T^{\ast})$, where $T^{\ast}$ is the maximum time of existence of (C.1), where we have used that for $|x-y|>1$ we have $\ln|x-y|\leqslant|x|+|y|$. The conclusion then follows from the bound on $I(t)$.

### v2

@@ C.1. No collision in finite time and global solutions for the second limiting system
Although no collisions in finite time is probably well-known, we provide the proof here since we did not find it in the literature. By Cauchy-Lipschitz theorem, there exists a unique local $C^{1}$ solution of (C.1).
As observed in [20], System (C.1) is a Hamiltonian system for the energy
$H_{0}:=\frac{1}{2}\sum_{i\neq j}\ln|q_{i}(t)-q_{j}(t)|+\frac{1}{4(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)^{2}.$
We can indeed check that this quantity is conserved.
Lemma C.1.
Let $t\mapsto(q_{1},\dots,q_{N})(t)$ a $C^{1}$ solution on $[0,T]$ of (C.1), then
$\frac{\mathrm{d}H_{0}}{\mathrm{d}t}=0.$
Proof.
We differentiate $H_{0}$ to obtain
$\frac{\mathrm{d}H_{0}}{\mathrm{d}t}=$ $\frac{1}{2}\sum_{i\neq j}\frac{q_{i}(t)-q_{j}(t)}{|q_{i}(t)-q_{j}(t)|^{2}}\cdot(\dot{q}_{i}(t)-\dot{q}_{j}(t))+\frac{1}{2(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)\dot{q}_{i,r}(t)$
(C.2) $=$ $\sum_{i\neq j}\frac{q_{i}(t)-q_{j}(t)}{|q_{i}(t)-q_{j}(t)|^{2}}\cdot\dot{q}_{i}(t)+\frac{1}{2(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)\dot{q}_{i,r}(t).$
By using (C.1), we compute the first right-hand-side sum
$\sum_{i\neq j}\frac{q_{i}-q_{j}}{|q_{i}-q_{j}|^{2}}\cdot\dot{q}_{i}=$ $\frac{\gamma}{2\pi}\sum_{i=1}^{N}\sum_{j\neq i}\sum_{k\neq i}\frac{q_{i}-q_{j}}{|q_{i}-q_{j}|^{2}}\cdot\frac{(q_{i}-q_{k})^{\perp}}{|q_{i}-q_{k}|^{2}}-\frac{\gamma}{4\pi(r^{*})^{2}}\sum_{i=1}^{N}\sum_{j\neq i}\frac{q_{i}-q_{j}}{|q_{i}-q_{j}|^{2}}\cdot{q}_{i,r}e_{z}$
$=$ $\frac{-\gamma}{4\pi(r^{*})^{2}}\sum_{i=1}^{N}\sum_{j\neq i}q_{i,r}\frac{q_{i,z}-q_{j,z}}{|q_{i}-q_{j}|^{2}}$
by inverting $k$ and $j$ in the first term.
The last sum in (C.2) is computed in the same way:
$\frac{1}{2(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}\dot{q}_{i,r}=$ $\frac{1}{2(r^{\ast})^{2}}\frac{\gamma}{2\pi}\sum_{i=1}^{N}\sum_{j\neq i}q_{i,r}e_{r}\cdot\frac{(q_{i}-q_{j})^{\perp}}{|q_{i}-q_{j}|^{2}}=\frac{\gamma}{4\pi(r^{\ast})^{2}}\sum_{i=1}^{N}\sum_{j\neq i}q_{i,r}\frac{q_{i,z}-q_{j,z}}{|q_{i}-q_{j}|^{2}}$
where we recall our convention $e_{z}=(1,0)$, $e_{r}=(0,1)$ and $a^{\perp}=(-a_{r},a_{z})$. This completes the proof of the conservation of the energy $H_{0}$. ∎
@@
$I(t):=\frac{1}{2}\sum_{i=1}^{N}|q_{i}(t)|^{2}$
satisfies
$\frac{\mathrm{d}}{\mathrm{d}t}I(t)$ $=\frac{\gamma}{2\pi}\sum_{i\neq j}q_{i}(t)\cdot\frac{(q_{i}(t)-q_{j}(t))^{\perp}}{|q_{i}(t)-q_{j}(t)|^{2}}-\frac{\gamma}{4\pi(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)q_{i,z}(t)$
$=-\frac{\gamma}{2\pi}\sum_{i\neq j}\frac{q_{i}(t)\cdot q_{j}^{\perp}(t)}{|q_{i}(t)-q_{j}(t)|^{2}}-\frac{\gamma}{4\pi(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)q_{i,z}(t)$
$=-\frac{\gamma}{4\pi(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)q_{i,z}(t)\leqslant\frac{\gamma}{8\pi(r^{\ast})^{2}}I(t),$
where we have exchanged $i$ and $j$ in the first sum, hence
$I(t)\leqslant I(0)e^{\gamma t/(8\pi r^{\ast 2})}.$
As a consequence, using $\ln|x-y|\leqslant|x-y|\leqslant|x|+|y|$, we get
$\sum_{i\neq j||q_{i}-q_{j}|<1}|\ln|q_{i}(t)-q_{j}(t)||$ $=\sum_{i\neq j||q_{i}-q_{j}|>1}\ln|q_{i}(t)-q_{j}(t)|-2H_{0}+\frac{1}{2(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)^{2}$
$\leqslant 2N\sum_{i=1}^{N}|q_{i}(t)|-2H_{0}+\frac{1}{2(r^{\ast})^{2}}I(t)$
$\leqslant 2N\sqrt{N}\sqrt{I(t)}-2H_{0}+\frac{1}{2(r^{\ast})^{2}}I(t),$
for all $t\in[0,T^{\ast})$, where $T^{\ast}$ is the maximum time of existence of (C.1). The conclusion of non-collision in finite time then follows from the bound on $I(t)$. By the estimate on $I$, we also know that the particles can not escape to infinity in finite time, which implies that the unique $C^{1}$-solution of (C.1) is global.
C.2. Difference between the two dynamics
We provide numerical evidence in the case $N=2$ that the dynamics of $Y_{i}^{\varepsilon}(t)$ as defined in (1.7) differs from the limit dynamics for $\widetilde{Y}_{i}$ obeying (1.8). For convenience, we choose $X^{*}=(X_{1}^{\varepsilon}(0)+X_{2}^{\varepsilon}(0))/2$ so that $Y_{1}(0)+Y_{2}(0)=0$. Let us recall that we proved that as $\varepsilon\to 0$, then $Y_{i}^{\varepsilon}(t)\to\widetilde{Y}_{i}(t)$ uniformly in $t\in[0,T_{0}]$, see Lemma 5.3. However, for a given $\varepsilon>0$, the respective dynamics of the points $Y_{i}^{\varepsilon}(t)$ and $\widetilde{Y}_{i}(t)$ have different properties.
First, we observe that for the second system, the center of mass $(\widetilde{Y}_{1}+\widetilde{Y}_{2})/2$ does not move:
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{\widetilde{Y}_{1}+\widetilde{Y}_{2}}{2}=-\frac{\gamma}{4\pi(r^{*})^{2}}\frac{\widetilde{Y}_{1,r}+\widetilde{Y}_{2,r}}{2}e_{z},$
which implies that $\widetilde{Y}_{1,r}+\widetilde{Y}_{2,r}$ remains constant. As we have assumed here that this is initially zero, the previous equality also gives that the vertical component is at rest, so in the end, the $(\widetilde{Y}_{1}+\widetilde{Y}_{2})/2$ is constant. Then, by showing that every level set of the associated Hamiltonian is bounded, it is possible to deduce from this that the trajectories of the system $(\widetilde{Y}_{1},\widetilde{Y}_{2})$ are periodic in time for every initial data, see [20].
This property on the center of mass is false for the system $(Y_{1}^{\varepsilon},Y_{2}^{\varepsilon})$ as can be seen by the following computation. By taking the derivative in time of the center of mass, we get that
(C.3) $\frac{\mathrm{d}}{\mathrm{d}t}\frac{Y_{1}^{\varepsilon}+Y_{2}^{\varepsilon}}{2}=\sqrt{|\ln\varepsilon|}\frac{\gamma}{8\pi r^{*}}\left(\frac{1}{1+\frac{Y_{1,r}^{\varepsilon}}{r^{*}\sqrt{|\ln\varepsilon|}}}+\frac{1}{1+\frac{Y_{2,r}^{\varepsilon}}{r^{*}\sqrt{|\ln\varepsilon|}}}-2\right)e_{z}$
which has again no radial part, but cannot vanish at all times. Moreover, by noticing that the radial components are bounded and opposite at every time, we perform a Taylor expansion of (C.3) up to the second order to get that
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{Y_{1}^{\varepsilon}+Y_{2}^{\varepsilon}}{2}$ $=\sqrt{|\ln\varepsilon|}\frac{\gamma}{8\pi r^{*}}\left(-\frac{Y_{1,r}^{\varepsilon}+Y_{2,r}^{\varepsilon}}{r^{*}\sqrt{|\ln\varepsilon|}}+\frac{(Y_{1,r}^{\varepsilon})^{2}+(Y_{2,r}^{\varepsilon})^{2}}{(r^{*})^{2}|\ln\varepsilon|}+\mathcal{O}\left(\frac{1}{|\ln\varepsilon|^{3/2}}\right)\right)e_{z}$
$=\frac{\gamma}{8\pi}\frac{(Y_{1,r}^{\varepsilon})^{2}+(Y_{2,r}^{\varepsilon})^{2}}{(r^{*})^{3}\sqrt{|\ln\varepsilon|}}e_{z}+\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right).$
This quantity being non-negative, we expect the dynamics of $(Y_{1}^{\varepsilon},Y_{2}^{\varepsilon})$ to display a non-linear shift towards positive $z$. This fact is observed numerically as shown in Figure 3.
[…]
(a) Trajectories of $Y_{1}^{\varepsilon}(t)$ and $Y_{2}^{\varepsilon}(t)$
(b) Trajectory of $Y_{1}^{\varepsilon}(t)$
(c) Time evolution of $(Y_{1,z}^{\varepsilon}(t)+Y_{2,z}^{\varepsilon}(t))/2$
(d) Vertical velocity of $(Y_{1}^{\varepsilon}(t)+Y_{2}^{\varepsilon}(t))/2$
Figure 3. Numerical computation of the motion in ${\mathbb{R}}^{2}$ of $Y_{1}^{\varepsilon}(t)$ and $Y_{2}^{\varepsilon}$ during a full rotation (A), and the same motion of $Y_{1}^{\varepsilon}(t)$ during three full rotations (B). The shift towards positive $z$ is seen in (C), with a non-trivial velocity (D). The parameters used are $\varepsilon=0.01$, $\gamma=2\pi$, $r^{*}=1$ and $Y_{1}(0)=(0,1.2)$.
To understand more in detail the dynamics, one could introduce the corrected positions made such that the center of mass of the configuration is preserved, namely the points
$\hat{Y}_{i}^{\varepsilon}(t)=Y_{i}^{\varepsilon}(t)-\frac{Y_{1}^{\varepsilon}(t)+Y_{2}^{\varepsilon}(t)}{2},$
which satisfy by construction that $\hat{Y}_{1}^{\varepsilon}=-\hat{Y}_{2}^{\varepsilon}$, meaning that the dynamics reduces to the study of a single point in ${\mathbb{R}}^{2}$, allowing to obtain the phase portrait type Figure 4, displaying different trajectories of $\hat{Y}_{1}^{\varepsilon}(t)$ (and thus of $\hat{Y}_{2}^{\varepsilon}(t)$) for various initial data. In particular, we recover a periodic motion, with a phase portrait very similar to that of the $\widetilde{Y}_{i}$, see [20]. In conclusion, we claim that the dynamics of $(Y_{1}^{\varepsilon},Y_{2}^{\varepsilon})$ differs from the dynamics of $(\widetilde{Y}_{1},\widetilde{Y}_{2})$ essentially by a non-trivial shift towards the positive $z$.
Figure 4. Phase portrait of the centered dynamics $(\hat{Y}_{1}^{\varepsilon},\hat{Y}_{2}^{\varepsilon})$.